KPSESTAT(1) KPSESTAT(1) NAME kpsestat−computeoctalmodefrommodeofexistingfile SYNOPSIS kpsestatmodefile DESCRIPTION kpsestatprintstheoctalpermissionoffilemodifiedaccordingtomodeonstandardoutput.The modeparameteracceptsasubsetofthesymbolicpermissionsacceptedbychmod(1).Use=as themodetoobtaintheunchangedpermissions. OPTIONS kpsestatacceptsthefollowingadditionaloptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. SEEALSO chmod(1). Kpathsea7.32x 4January1998 1